Tohban Report 2018-04-18
Jump to navigation
Jump to search
Tohban Reports | |
---|---|
Start Date: | 11 April 2018 |
End Date: | 18 April 2018 |
Tohban: | Jim McTiernan |
Tohban email: | jimm@ssl.berkeley.edu |
Next Tohban: | TBD |
List all reports |
Solar Activity
Quiet...
Some GOES activity, No RHESSI flares to count.
Memory Management
Detectors were turned off for annealing at 21:35 on April 12th; Decimation was active/vigorous at the time. The SSR read pointer was moved to 135000000 on 17-April, to reduce the SSR fill to 7.5%. This resulted in a vc3 data gap from 11-apr-2018 01:50 until 16-apr-2018 02:30. Currently the SSR fill is approximately 10%.
Spacecraft Status
Detectors are off:
http://sprg.ssl.berkeley.edu/~ayshih/soh/latest_temperatures.pdf
for a summary plot of recent temperatures, efficiencies and accelerometer values.
Data Gaps
DATA GAPS FOR 2018-04-11 To 2018-04-18
TIME RANGE: 2018-04-11T00:00:00.000 -- 2018-04-18T00:00:00.000 NO GAPS IN APP_ID = 1 (VC1-SOH) NO GAPS IN APP_ID = 154 (VC1-PMTRAS) GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 3 GAP START TIME GAP END TIME GAP (SEC) 2018-04-11T00:10:00.000 -- 2018-04-11T00:15:00.000 300.00000 2018-04-11T01:50:00.000 -- 2018-04-16T02:30:00.000 434400.00 2018-04-17T05:05:00.000 -- 2018-04-18T00:00:00.000 68100.000
Detector issues
Detectors are off.
Spacecraft Management
Decimation | Active/Vigorous |
HLAT Decimation | Rear decimation |
Night time data (fronts) | plus/minus 4 minutes from daylight |
Night time data (rears) | plus/minus 4 minutes from daylight |
Require extra passes? | No |
Requirement for moving pointer? | See above |
Attenuator operation | Nominal |
Ref. deadtime for Att.thresholds | 35.x% |
Detector problems? | See above |